OpenStudy (anonymous):

1. Find the slope of line passing through the points (5, 9) and (5, -2) A. -11 B. 0 C. 11 D. undefined 2. Find the slope of the line passing through the points (-2, 9) and (5, 3). A. -6/7 B. 4 C. -7/6 D. 1/4 3. Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (10, -3) with a slope of -3/2. A. -13 B. 4 C. 13 D. -12 4. In a linear equation, the independent variable increases at a constant rate while the dependent variable decreases at a constant rate. The slope of this line is A. zero B. negative C. positive D. undefined

OpenStudy (unklerhaukus):

1. D -undefined both points have the same x co-ordinate , thus the slope is vertical and undefined

OpenStudy (anonymous):

r u sure?

OpenStudy (unklerhaukus):

slope of a line is rise/run, rise is -11 run is 0 -11/0 is Undefined

OpenStudy (anonymous):

oh okay. Thank you for tht explanation.

OpenStudy (anonymous):

The slope of two points (x1,y1) and (x2,y2) is given by the formula m=(y2-y1)/(x2-x1) Plug your points into the formula and you get m=(3-9)/(5-(-2))=(-6)/(7) So wouldn't it be -6/7

OpenStudy (unklerhaukus):

so for 2. the slope from (-2,9) to (5,3) is rise over run which is -6/7 = -6/7 which is A.

OpenStudy (unklerhaukus):

yeah that is right

OpenStudy (anonymous):

okay i thought so

OpenStudy (unklerhaukus):

can you do the others by your self ?

OpenStudy (anonymous):

well yea i already done them i was just having you guys check it.

OpenStudy (anonymous):

Can you help me with 3?